40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 20 19:09:48 2015, ericq, Update, ASC, QPD frontend code unified 
    Reply  Mon Feb 23 08:06:59 2015, rana, Update, ASC, QPD frontend code unified 
    Reply  Thu Mar 5 20:44:30 2015, Jenne, Update, SUS, damprestore script updated 
    Reply  Mon Mar 9 18:10:59 2015, Jenne, Update, ASC, Oplev filters re-copied 
    Reply  Sat Mar 14 00:12:18 2015, rana, Update, SUS, Oplevs huh? burp.png
       Reply  Mon Mar 23 13:27:33 2015, ericq, Update, SUS, ITMX oplev quadrant gains unbalanced olsum.png
          Reply  Fri Apr 3 14:57:38 2015, manasa, Update, SUS, BS oplev BS_oplev_Apr3.png
             Reply  Fri Apr 3 15:15:55 2015, Steve, Update, SUS, BS oplev 
                Reply  Fri Apr 3 19:35:14 2015, Jenne, Update, SUS, BS oplev centered 
Message ID: 11199     Entry time: Fri Apr 3 14:57:38 2015     In reply to: 11160     Reply to this: 11200
Author: manasa 
Type: Update 
Category: SUS 
Subject: BS oplev 

The BS oplev has been misbehaving and kicking the optic from time to time since noon. The kicks are not strong enough to trip the watchdogs (current watchdog max counts for the sensors is 135).

I took a look at the spectrum of the BS oplev error in pit and yaw with both loops enabled while the optic was stable. There is nothing alarmingly big except for some additional noise above 4Hz.

I have turned the BS oplev servo OFF for now.

Attachment 1: BS_oplev_Apr3.png  83 kB  | Hide | Hide all
BS_oplev_Apr3.png
ELOG V3.1.3-